OpenStudy (anonymous):

Find (a) the equation of a line that is parallel to the given line and includes the given point, and (b) the equation of a line that is perpendicular to the given line through the given point. Write both answers in slope-intercept form. (c) Graph both of these lines on the same axes and submit your graph to the Dropbox. y = 2x 6, (-1, -2)

OpenStudy (campbell_st):

is the line y = 2x + 6...?

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

(a) the gradient of the parallel line is 2, the same are the given line then use the point slope formula to find the equation \[y - y _{1} = m(x -x _{1})\] simplify and write in point slope form... (x\1 an y\1) come from the given point and m = 2 (b) the perpendicular has a gradient of -1/m . The value of m was found in part a. use the formula and point again...

OpenStudy (anonymous):

ok